OpenStudy (anonymous):

y=-2/5x+1 a. 2x+5y=5 b. -5x=2y+5 c. x=-2/5y+1 d. -5y=2x+5

OpenStudy (anonymous):

What exactly is the question/instructions?

OpenStudy (anonymous):

Write the equation in standard form:

OpenStudy (compassionate):

Standard form is Ax + By = C

OpenStudy (anonymous):

y = (2/5)x + 1 Multiply both sides by 5. 5y = 2x + 5 Convert that form into Standard form. 2x + 5y = 5

OpenStudy (anonymous):

So I believe A would be the answer

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Sure thing